Circular No. 7794

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVAE IN M31 Independent discoveries of three apparent novae in M31 have been

received from M. Fiaschi, F. Di Mille, and R. Cariolato, Astronomical Observatory ‘G. Colombo’, Padua, and from B. Swift and W. D. Li, University of California at Berkeley. The Padua images were taken with a 0.41-m reflector through Hα and R filters, while the unfiltered Berkeley images were taken in the course of the LOTOSS (cf. IAUC 7514) with the Katzman Automatic Imaging Telescope (KAIT). The table gives the Padua positions, offsets, and discovery magnitudes:

2002 UT  α_{2000}  δ_{2000}  H_{α} R Offset
Jan. 13.77 0^{h}42^{m}33.^{s}79 +41^{o}18′23′′.7 15.7 18.4 157′′ W, 135′′ N
Jan. 13.77 0 42 41.38 +41 16 24.6 14.6 17.3 43′′ W, 16′′ N
Jan. 13.77 0 42 52.83 +41 15 10.6 15.9 18.3 128′′ E, 58′′ S

On Jan. 14.78 UT, the Hα magnitudes for the respective apparent novae were 16.0, 14.7, and 16.1; nothing was visible at these locations in Padua Hα images (FWHM 6.3 nm) taken on 2001 Sept. 15 (limiting mag 17.5). KAIT position end figures (with offsets) and magnitudes for each of the three apparent supernovae on 2002 Jan. 12.1 and 13.1 are: 33^{s}.94, 24′′.0 (118′′ west, 135′′ north), ∼ 17.5; 41^{s}.26, 24′′.2 (33′′ west, 16′′ north), ∼ 18.0; 52^{s}.75, 10′′.8 (97′′ east, 58′′ south) ∼ 17.5. None of the three new stars was visible on KAIT images taken on 2001 Dec. 19.2 (limiting mag ∼ 19.0).

SATURN VIII (IAPETUS) G. Black, National Radio Astronomy Observatory; D. Campbell, L.

Carter, Cornell University/NAIC; and S. Ostro, Jet Propulsion Laboratory, announce the detection of radar echoes from Saturn VIII (Iapetus) on Jan. 8–10 using the Arecibo Observatory’s 13-cm-wavelength radar: “The observations were of the brighter trailing hemisphere and centered at a satellite longitude of 260^{o} west. The radar scattering properties of Iapetus are very different from those of the icy Galilean satellites, with the cross section significantly lower and the circular polarization ratio < 1.”

COMET P/2001 Y1 (HELIN-LAWRENCE) Further to IAUC 7790, T. Oribe reports that his CCD images taken

on 2001 Dec. 24.86 UT show a 0′.15 coma and an 8′′ tail in p.a. 295^{o}.

2002 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green